# Search Timelines: Visualizing Search History to Enable Cross-Session Exploratory Search

### v1

@@ Abstract
Purpose: The timespan over which exploratory searching can occur, as well as the scope and volume of the search activities undertaken, can make it difficult for searchers to remember key details about their search activities. These difficulties are present both in the midst of searching as well as when resuming a search that spans multiple sessions. In this paper, we present a search interface designed to support cross-session exploratory search in a public digital library context. Methods: Search Timelines provides a visualization of current and past search activities via a dynamic timeline of the search activity (queries and saved resources). This timeline is presented at two levels of detail. An overview timeline is provided alongside the search results in a typical search engine results page design. A detailed timeline is provided in the workspace, where searchers can review the history of their search activities and their saved resources. A controlled laboratory study was conducted to compare this approach to a baseline interface modelled after a typical public digital library search/workspace interface. Results: Participants who used Search Timelines reported higher levels of user engagement, usability, and perceived knowledge gain, during an initial search session and when resuming the search after a 7-8 day interval. This came at the expense of the searchers taking more time to complete the search task, which we view as positive evidence of engagement in cross-session exploratory search processes. Conclusion: Search Timelines serves as an example of how lightweight visualization approaches can be used to enhance typical search interface designs to support exploratory search. The results highlight the value of providing persistent representations of past search activities within the search interface.
@@
Competing Interest: This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), through the Alliance Grant (ALLRP 558319-20) held by the first author, in collaboration with the fourth and fifth authors, the Regina Public Library, and the Saskatchewan Information and Library Services Consortium.
@@ 1 Introduction
A public digital library is an online platform that provides access to a wide range of library resources (including books, multimedia materials, and other educational content) to the general public [20, 24]. Public digital libraries allow users to access and search for library holdings without the constraints of traditional library hours or physical locations. When searching for known or easily described items, the simple search interfaces provided by public digital libraries can help library patrons find the resources they are seeking. However, such simple interfaces provide limited support to searchers when their tasks become complex [15].
@@
To resolve complex search tasks, searchers may undertake searches that span multiple sessions over an extended period of time [2, 21]. To resume and reacquaint themselves with the previous search activities, searchers need explicit support from the search interface [28, 21, 22]. Unfortunately, while performing cross-session searching, searchers are not well supported by the majority of public digital library search interfaces, as little information is available to help searchers reacquaint themselves with their past search activities beyond a simple list of saved resources in a workspace. Because of this, searchers are forced to reacquaint themselves and resume the past search tasks on their own, and must come up with individual strategies for carrying out cross-session searches [21, 22].
@@
The remainder of this paper is organized as follows. A literature review is provided in Section 2, establishing the foundations for this research. A detailed description of the design of Search Timelines is provided in Section 3. The methods for evaluating this approach in comparison to a typical baseline system are provided in Section 4. The results of this evaluation are provided in Section 5, followed by a discussion on how these findings relate to previous work and what new knowledge has been discovered in Section 6. The paper concludes with a summary of the primary contributions and an overview of future work in Section 7.
@@ 2 Literature Review
Of particular relevance to this research are the timeline used in LogCanvas [42] and the hierarchical representation used in SearchBar [28]. Both approaches provide mechanisms for evaluating past search activities, but go about it in different ways. LogCanvas uses a vertical timeline to represent search sessions. For each search session, the start and end time are specified, along with a list of all queries that were issued during the session. This approach allows a searcher to see when queries were previously issued, but without an ability to assess their success in surfacing relevant search results. SearchBar uses a collapsible tree structure representation of search topic, queries issued, and search results viewed. This approach allows a searcher to see which queries lead to relevant search results (those the searcher chose to view or mark as useful), but without an ability to see when the queries were issued or when individual search results were viewed.
@@ 3.1 Search Topic Management
A fundamental principle in search topic management is that each search topic must preserve a separate history and list of saved search results. We have implemented a simple approach for such workspaces that follows a pattern that has been shown to be effective in other approaches [10]. When beginning a new session, the searcher’s first query becomes the topic title for the workspace. If the searcher finds at some point that this is no longer an accurate description of the topic, it can be edited from within the workspace.
Using the concept of an ongoing (current) search topic, all user actions including issuing new queries, saving search results, and removing saved search results, occur within the context of the ongoing search topic’s workspace. This ongoing topic also determines what information to display in the timeline visualizations. A list of all available workspaces is also provided, enabling the searcher to easily switch between search topics, as well as create new ones.
@@
Figure 1: The SERP view of Search Timelines, displaying results for the query, “indigenous people”. Two documents from this query have already been added to the ongoing search topic, “First Nations”, and can be seen in the overview timeline to the right.
@@ 3.2 Workspace and Detailed Timeline
The information saved to the workspace is organized as a detailed timeline, which merges the design of the overview timeline with the design of the search results cards from the SERP (see Figure 2). The mechanism for linking the icons of the saved search results back to their source queries is the same between both timelines, although here the icons are larger and the lines connecting them to the queries are thicker and darker. Rather than only listing the title of the saved search result (as is done in the overview timeline), the complete search result card is used.
Figure 2: The workspace view of Search Timelines, with “First Nations” as the displayed topic and “Canadian Landscape” as the ongoing topic.
The workspace can be used in the context of an ongoing search task, providing searchers with the ability to assess the details of the saved search results. This an important feature in the context of an exploratory search process since the searchers’ understanding of their topics will change as they discover new information and learn about the topic [40]. Any previously saved search results that are determined to no longer be relevant can be removed from the workspace by clicking on the “remove” button. In order to preserve the timeline of the search activity, this does not completely delete the item, but instead uses a strike-through on the title and reduces the saturation of the font and icons to minimize the visual presence of these search results. This occurs for both the detailed timeline and the overview timeline.
@@
Figure 3: The Baseline interface displaying results for the query, “First Nations”. Resources can be saved to the workspace with the “add to saved documents button”. The workspace has a similar visual representation of the resources that have been saved.
@@ 4.1 Methodology
A between-subjects controlled laboratory study design [18] was used to evaluate the search interface, with a longitudinal component spanning two sessions. The inclusion of this longitudinal aspect was essential to capture the dynamics of cross-session searching. Moreover, this study allowed for an evaluation of how search behaviour might be influenced by extended gaps between sessions and how searchers approached the resumption of a previously unfinished task.
@@ 4.2 Study Procedures
Using a between-subjects design, each participant followed the same study procedures, with the only difference being that half the participants were given Search Timelines to use for the search task while the other half were given the Baseline. Details of the study procedures are illustrated in Figure 4.
Of particular importance in this study design is the creation of an environment where the task cannot be completed in a single search session. To do so, we limited the initial search to a maximum of 15 minutes and scheduled the second search session to occur 7-8 days later. The second search session had no time limit. The rest of the study followed a typical interactive information retrieval user study pattern. The entire process was reviewed and approved by the Research Ethics Board at our university.
@@ 4.3 Simulated Search Task
Search Scenario: Suppose you are a high school student and taking a class on History. You have been given a class assignment to write a brief report on the French Revolution. The French Revolution was a period of political and social upheaval in France from 1789 to 1799. The Revolution was triggered by a combination of factors, including a financial crisis, food shortages, and widespread dissatisfaction with the absolute monarchy and the privileges of the aristocracy. Initial Search Instructions: Imagine yourself in a history class. Your teacher has given you some time at the end of class to start your research for the assignment on the French Revolution. Since the class will be over in 15 minutes, you have decided that you would like to spend your time learning about the variety of information that is available about this topic. Your teacher has told you that you should not use web pages as your source of information, so you decide to search for books in the public library and save any you think will help you with your assignment. You have given yourself the goal of learning as much as you can about the topic, and to not make any decisions on the specific focus of your assignment. Resumed Search Instructions: About a week ago, you spent fifteen minutes in your history class looking for some resources to get you started with your assignment on the French Revolution. The French Revolution was a period of political and social upheaval in France from 1789 to 1799. The Revolution was triggered by a combination of factors, including a financial crisis, food shortages, and widespread dissatisfaction with the absolute monarchy and the privileges of the aristocracy. Imagine you are now at home and have plenty of time to continue your previously unfinished search task of finding relevant resources for your assignment. Since your prior search activities focused on learning about the topic, your goal now is to pick up where you left off, and then focus on a specific aspect of the French Revolution so that your paper can have a narrow focus. Continue searching within the public library, save what you like, and delete those that are no longer relevant.
@@ 4.4 Participants
32 participants were recruited from the undergraduate population at our university, with the first half assigned the Search Timelines interface and the second half assigned the Baseline. One participant assigned to the Baseline condition dropped out after the first session; the data for that participant was discarded and another participant was recruited.
@@
Since we used a convenience sample of participants (students recruited broadly from within our University student population), we took care to situate them within a context where they would need to search within the public library, rather than the university library. We did this by asking the participants to imagine being a high school student, instead of a university student.
Although the participants themselves were no longer high schools students (as identified in the simulated work task), it was necessary to situate them in this search context because our data source was a public library and not an academic (university) library. None of the participants expressed concerns with having to image being a high school student.
@@ 4.5 Analysis
With one independent variable, the data was analyzed using statistical comparisons between the participants who used Search Timelines versus those who used the Baseline. The survey data was aggregated per concept: aspects of and overall user engagement (RQ1a and RQ1b), aspects of usability (RQ2a and RQ2b), and perceived knowledge gain (RQ3a and RQ3b). The time to task completion data was used as-is (RQ4a/b). Shapiro-Wilk tests were used to assess whether the data was normally distributed.
For each of the statistical comparisons of the survey data, at least one set of data was found to not be statistically significant. As a result, the non-parametric Mann-Whitney U test was used, along with Cohen’s d for effect size measurements. Of note, when there are similar data across different measures, it is possible for the test statistic W to be the same value (in our case, 256).
For the statistical comparison of the time to task completion data, the distribution for each condition was normal. As a result, ANOVA tests were used along with eta-squared for effect size measurements.
@@ 5.1 RQ1a: User Engagement During the Initial Search
Participants using Search Timelines indicated higher levels of user engagement compared to those who used the Baseline, which were found to be statistically significant with large effect sizes. The aggregated data for each condition, along with the statistical analysis are provided in Table 1 for each aspect of user engagement, as well as overall.
@@ 5.2 RQ1b: User Engagement During the Resumed Session
When resuming the search task, participants who used Search Timelines reported higher user engagement scores compared to those using the Baseline, mirroring the results observed during the initial session. These findings were statistically significant with large effect sizes (see Table 2).
@@ 5.3 RQ2a: Usability During the Initial Search
Participants who used Search Timelines also reported higher usability scores than those who used the Baseline, measured through aspects of ease-of-use, usefulness, and satisfaction. As shown in Table 3, these findings were statistically significant with large effect sizes.
@@
Ease-of-use 2.67 (0.472) 4.95 (0.101) W = 256, p <$\mathbf{10^{-6}}$ d = 6.686
Usefulness 2.47 (0.386) 4.94 (0.144) W = 256, p <$\mathbf{10^{-6}}$ d = 8.473
Satisfaction 2.45 (0.306) 4.92 (0.198) W = 256, p <$\mathbf{10^{-6}}$ d = 9.580
@@
Ease-of-use 2.59 (0.364) 4.97 (0.085) W = 256, p <$\mathbf{10^{-6}}$ d = 8.990
Usefulness 2.20 (0.332) 4.95 (0.188) W = 256, p <$\mathbf{10^{-6}}$ d = 10.202
Satisfaction 2.58 (0.285) 4.92 (0.176) W = 256, p <$\mathbf{10^{-6}}$ d = 9.905

### v2

@@ Abstract
Purpose: The timespan over which exploratory searching can occur, as well as the scope and volume of the search activities undertaken, can make it difficult for searchers to remember key details about their search activities. These difficulties are present both in the midst of searching as well as when resuming a search that spans multiple sessions. In this paper, we present a search interface design and prototype implementation to support cross-session exploratory search in a public digital library context. Methods: Search Timelines provides a visualization of current and past search activities via a dynamic timeline of the search activity (queries and saved resources). This timeline is presented at two levels of detail. An Overview Timeline is provided alongside the search results in a typical search engine results page design. A Detailed Timeline is provided in the workspace, where searchers can review the history of their search activities and their saved resources. A controlled laboratory study (n=32) was conducted to compare this approach to a baseline interface modelled after a typical public digital library search/workspace interface. Results: Participants who used Search Timelines reported higher levels of user engagement, usability, and perceived knowledge gain, during an initial search session and when resuming the search after a 7-8 day interval. This came at the expense of the searchers taking more time to complete the search task, which we view as positive evidence of engagement in cross-session exploratory search processes. Conclusion: Search Timelines serves as an example of how lightweight visualization approaches can be used to enhance typical search interface designs to support exploratory search. The results highlight the value of providing persistent representations of past search activities within the search interface.
@@
[…]
Statements and Declarations Competing Interest: This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), through the Alliance Grant (ALLRP 558319-20) held by the first author, in collaboration with the fourth and fifth authors, the Regina Public Library, and the Saskatchewan Information and Library Services Consortium.
@@ 1 Introduction
A public digital library is an online platform that provides access to a wide range of library resources (including books, multimedia materials, and other educational content) to the general public [21, 25]. They allow library patrons to search for and access library holdings without the constraints of library hours or physical locations. The physical libraries often have multiple branches within a municipality, and may be part of a regional consortium to provide economies of scale to their digital library services.
The search interfaces provided by the public digital library follow common design patterns that are prevalent across many digital library contexts (e.g., a query box and a list of search results). When searching for known or easily described items, the simple search interfaces provided by public digital libraries can help library patrons find the resources they are seeking. However, such simple interfaces provide limited support to searchers when their tasks become complex [16].
@@
To resolve complex search tasks, searchers may undertake searches that span multiple sessions over an extended period of time [2, 22]. That is, they may start a search session and perform some initial (exploratory browsing) searching to familiarize themselves with what is available. As they discover new information, they may wish to pause their searching to give themselves an opportunity to think about what was found. Or they may have time constraints that keep them from finishing the task.
To resume and reacquaint themselves with the previous search activities, searchers need explicit support from the search interface [29, 22, 23]. Unfortunately, while performing cross-session searching, searchers are not well supported by the majority of public digital library search interfaces, as little information is available to help searchers reacquaint themselves with their past search activities beyond a simple list of saved resources in a workspace. Because of this, searchers are forced to reacquaint themselves and resume the past search tasks on their own, and must come up with individual strategies for carrying out cross-session searches [22, 23].
@@
The remainder of this paper is organized as follows. A literature review is provided in Section 2, establishing the foundations for this research. A detailed description of the design of Search Timelines is provided in Section 3. The methods for evaluating this approach in comparison to a typical baseline system are provided in Section 4. The results of the evaluation are provided in Section 5, followed by a discussion on how these findings relate to previous work and what new knowledge has been discovered in Section 6. The paper concludes with a summary of the primary contributions and an overview of future work in Section 7.
@@ 2 Literature Review
Of particular relevance to this research are the timeline used in LogCanvas [43] and the hierarchical representation used in SearchBar [29]. Both approaches provide mechanisms for evaluating past search activities, but go about it in different ways. LogCanvas uses a vertical timeline to represent search sessions. For each search session, the start and end time are specified, along with a list of all queries that were issued during the session. This approach allows a searcher to see when queries were previously issued, but without an ability to assess their success in surfacing relevant search results. SearchBar uses a collapsible tree structure representation of search topics, queries issued, and search results viewed. This approach allows a searcher to see which queries lead to relevant search results (those the searcher chose to view or mark as useful), but without an ability to see when the queries were issued or when individual search results were viewed.
@@ 3.1 Search Topic Management
Search topics represent the different search interests a user wishes to keep separate from one another. A fundamental principle in search topic management is that each search topic must preserve a separate history and list of saved search results. To support this, we have implemented a simple workspace approach that follows a pattern that has been shown to be effective in other approaches [11]. When beginning a new session, the searcher’s first query becomes the topic title for the workspace. If the searcher finds at some point that this is no longer an accurate description of the topic, it can be edited from within the workspace.
Using the concept of an ongoing (current) search topic, all user actions including issuing new queries, saving search results, and removing saved search results, occur within the context of the ongoing search topic’s workspace. This ongoing topic also determines what information to display in the timeline visualizations. A list of all available workspaces is also provided, enabling the searcher to easily switch between search topics, as well as create new ones
3.2 SERP and Overview Timeline
@@
Figure 1: The SERP view of Search Timelines, displaying results for the query, “indigenous people”. Two documents from this query have already been added to the ongoing search topic, “First Nations”, and can be seen in the Overview Timeline to the right. The Overview Timeline also shows information from two earlier search sessions.
@@
One potential limitation of the Overview Timeline is that it could grow very large for search tasks that extend over long periods of time and include many queries and saved resources. Although the solutions to this problem are beyond the scope of this research, it could be addressed by limiting the number of sessions shown or providing an interactive collapsing and expanding of the sessions with the older sessions being collapsed by default.
@@
Figure 2: The workspace view of Search Timelines, with “First Nations” selected as the displayed topic along with the Detailed Timeline of the search activities for this topic. The most recent search session is shown, including the query and saved resources.
@@ 3.3 Workspace and Detailed Timeline
Information saved to the workspace is organized as a Detailed Timeline, which merges the design of the overview timeline with the design of the search results cards from the SERP (see Figure 2). The mechanism for linking the icons of the saved search results back to their source queries is the same between both timelines, although here the icons are larger and the lines connecting them to the queries are thicker and darker. Rather than only listing the title of the saved search result (as is done in the Overview Timeline), the complete search result card is used.
The workspace can be used in the context of an ongoing search topic, providing searchers with the ability to assess the details of the saved search results. This an important feature in the context of an exploratory search process since the searchers’ understanding of their topics will change as they discover new information and learn about the topic [41]. Any previously saved search results that are determined to no longer be relevant can be removed from the workspace by clicking on the “remove” button. In order to preserve the timeline of the search activity, this does not completely delete the item, but instead uses a strike-through on the title and reduces the saturation of the font and icons to minimize the visual presence of these search results. This occurs for both the Detailed Timeline and the Overview Timeline.
@@
Figure 3: The Baseline interface uses a common digital library search design pattern, displaying results for the query “First Nations”. Resources can be saved to the workspace with the “add to saved documents button”. The workspace uses a similar visual representation for the saved resources.
3.4 Prototype Implementation
The Search Timelines prototype was implemented using the MERN stack (MongoDB, Express, React, Node), and the MaterialUI library. Search queries are submitted to and processed by a RESTful API managed by a regional public library consortium, returning a ranked list of search results in JSON format. Each user’s workspaces, queries, and saved resources are stored in a centralized database.
@@ 4.1 Methodology
A between-subjects controlled laboratory study design [19] was used to evaluate the search interface, with a longitudinal component spanning two sessions. The inclusion of this longitudinal aspect was essential to capture the dynamics of cross-session searching. To avoid introducing a confounding factor related to topic/workspace switching, a between-subjects design allowed the use of a single search topic. Moreover, this study allowed for an evaluation of how search behaviour might be influenced by extended gaps between sessions and how searchers approached the resumption of a previously unfinished task.
@@ 4.2 Study Procedures
Each participant followed the same study procedures, with the only difference being that half the participants were given Search Timelines to use while conducting the search task, while the other half were given the Baseline. In order to help participants familiarize themselves with the assigned search interface, a pre-recorded training video was shown to the participants and they were then given an opportunity to use the interface to recreate the search scenario shown in the training video. A similar training session (video and experiential training) was provided for the second phase of the study, focusing on resuming the previous training scenario task. Other details of the study procedures are illustrated in Figure 4. The study was conducted online using Zoom, with screen sharing activated so that participant progress through the steps of the study could be monitored. A Qualtrics survey was used to provide links to the assigned interface, the training material, and to collect the survey data.
Of particular importance in this study design is the creation of an environment where the task cannot be completed in a single search session. To do so, we limited the initial search to a maximum of 15 minutes and scheduled the second search session to occur 7-8 days later. This time gap between sessions is consistent with other studies on cross-session search [29, 8, 11]. The second search session had no time limit. The rest of the study followed a typical interactive information retrieval user study pattern. The entire process was reviewed and approved by the Research Ethics Board at our university.
@@ 4.3 Simulated Search Task
Search Scenario: Suppose you are a high school student and taking a class on History. You have been given a class assignment to write a brief report on the French Revolution. The French Revolution was a period of political and social upheaval in France from 1789 to 1799. The Revolution was triggered by a combination of factors, including a financial crisis, food shortages, and widespread dissatisfaction with the absolute monarchy and the privileges of the aristocracy.
Initial Search Instructions: Imagine yourself in a history class. Your teacher has given you some time at the end of class to start your research for the assignment on the French Revolution.
Since the class will be over in 15 minutes, you have decided that you would like to spend your time learning about the variety of information that is available about this topic. Your teacher has told you that you should not use web pages as your source of information, so you decide to search for books in the public library and save any you think will help you with your assignment. You have given yourself the goal of learning as much as you can about the topic, and to not make any decisions on the specific focus of your assignment.
Resumed Search Instructions: About a week ago, you spent fifteen minutes in your history class looking for some resources to get you started with your assignment on the French Revolution. The French Revolution was a period of political and social upheaval in France from 1789 to 1799. The Revolution was triggered by a combination of factors, including a financial crisis, food shortages, and widespread dissatisfaction with the absolute monarchy and the privileges of the aristocracy.
Imagine you are now at home and have plenty of time to continue your previously unfinished search task of finding relevant resources for your assignment. Since your prior search activities focused on learning about the topic, your goal now is to pick up where you left off, and then focus on a specific aspect of the French Revolution so that your paper can have a narrow focus. Continue searching within the public library, save what you like, and delete those that are no longer relevant.
@@ 4.4 Participants
A convenience sample of 32 participants were recruited from the undergraduate population at our university, with the first half assigned the Search Timelines interface and the second half assigned the Baseline. One participant assigned to the Baseline condition dropped out after the first session; the data for that participant was discarded and another participant was recruited.
@@
A side-effect of using a convenience sample was that we needed to situate the participants within a context of searching within the public library, rather than the university library. This was necessary because our data source was connected to a public library not an academic (university) library. We did this by asking the participants to imagine being a high school student, instead of a university student. This could have been avoided by recruiting among high school students; however doing so would have introduced complications associated with obtaining school board and research ethics approval. Using convenience samples and simulated search tasks were deemed an acceptable compromise. Of note, none of the participants expressed concerns with having to imagine being a high school student.
@@ 4.5 Analysis
With one independent variable, the data was analyzed using statistical comparisons between the 16 participants who used Search Timelines versus the 16 who used the Baseline. The survey data was aggregated per concept: aspects of and overall user engagement (RQ1a and RQ1b), aspects of usability (RQ2a and RQ2b), and perceived knowledge gain (RQ3a and RQ3b). The time to task completion data was used as-is (RQ4a/b). Shapiro-Wilk tests were used to assess whether the data was normally distributed.
For each of the statistical comparisons of the survey data, at least one set of data was found to not be statistically significant. As a result, the non-parametric Mann-Whitney U test was used, along with Cohen’s d for effect size measurements. A p-value of $0.05$ was used to determine statistical significance. Of note, when there are large differences in the data being analyzed, it is possible for the test statistic W to be the same value across many of the measures (in our case, 256). The effect size is interpreted as small when $d=0.2$, medium when $d=0.5$, and large when $d\geq 0.8$.
For the statistical comparison of the time to task completion data, the distribution for each condition was normal. As a result, ANOVA tests were used along with eta-squared for effect size measurements. The effect size is interpreted as small when $\eta^{2}=0.01$, medium when $\eta^{2}=0.06$, and large when $\eta^{2}\geq 0.14$.
@@ 5.1 RQ1a: User Engagement During the Initial Search
Participants using Search Timelines indicated higher levels of user engagement compared to those who used the Baseline, which were found to be statistically significant with medium to large effect sizes. The aggregated data for each condition, along with the statistical analysis are provided in Table 1 for each aspect of user engagement, as well as overall.
@@ 5.2 RQ1b: User Engagement During the Resumed Session
When resuming the search task, participants who used Search Timelines reported higher user engagement scores compared to those using the Baseline, mirroring the results observed during the initial session. These findings were statistically significant with medium to large effect sizes (see Table 2).
@@ 5.3 RQ2a: Usability During the Initial Search
Participants who used Search Timelines also reported higher usability scores than those who used the Baseline, measured through aspects of ease-of-use, usefulness, and satisfaction. As shown in Table 3, these findings were statistically significant with medium to large effect sizes.
@@
Ease-of-use 2.67 (0.472) 4.95 (0.101) W = 256, p $\mathbf{10^{-6}}$ d = 6.686
Usefulness 2.47 (0.386) 4.94 (0.144) W = 256, p $\mathbf{10^{-6}}$ d = 8.473
Satisfaction 2.45 (0.306) 4.92 (0.198) W = 256, p $\mathbf{10^{-6}}$ d = 9.580
@@
Ease-of-use 2.59 (0.364) 4.97 (0.085) W = 256, p $\mathbf{10^{-6}}$ d = 8.990
Usefulness 2.20 (0.332) 4.95 (0.188) W = 256, p $\mathbf{10^{-6}}$ d = 10.202
Satisfaction 2.58 (0.285) 4.92 (0.176) W = 256, p $\mathbf{10^{-6}}$ d = 9.905